Wednesday Homework Problem 12.9 A Korean metamaterial formed of a lattice of conducting rectangles acheives an index of refraction of 38.6. What is the critical angle of the metamaterial in air?

(a) No critical angle exists for this material. (b) 1.1 ? (c) 1.5 ? (d) 2.1 ? . (e) Toby

Wednesday Homework Problem 12.10 If sunlight with intensity 1000W/m2 shines on an engineered material with index of refraction 50 at normal incidence, calculate the intensity of the light that is transmitted into the material.

(a) 39W/m2 (b) 961W/m2 (c) 77W/m2 (d) 923W/m2 (e) Toby.
Answer #1

1) c) 1.5 degrees

critical angle, theta_c = sin^-1(1/n)

= sin^-1(1/38.6)

= 1.5 degrees

2) c) 77 W/m^2

given, Io = 1000 W/m^2, n = 50

we know,
I = Io*4*n/(1 + n)^2

= 1000*4*50/(1 + 50)^2

= 77 W/m^2